A cup of coffee is $2.00 and the average variable expense per cup is $0.75. The average fixed expense per month is $2,000. An average of 2,500 cups are sold each month.

What is the CM Ratio for Coffee Klatch?

What is the Break-even Sales Dollars?

What is the Break-even sales in units?
Answer #1

Contribution Margin = Selling Price - Variable cost = 2.00-0.75 =1.25

CM ratio = contibution MArgin / Selling Price = 1.25/2.00 =62.50% (Answer)

Break Even Dollars = Fixed cost / CM ratio = 2,000/62.50% = $3,200 (Answer)

Break Even Units = fixed cost / CM Per Unit = 2000/1.25 =1,600 Units (Answer)